OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (anonymous):

@ursu27

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

@DLS

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (phi):

the ratio is L/W = 7/5 cross multiply to get 5L = 7W solve for L: divide both sides by 5 L =(7/5) W now write an equation for the perimeter using L and W can you do that ?

OpenStudy (phi):

can you write down the equation for the perimeter of a rectangle?

OpenStudy (phi):

see http://www.coolmath.com/reference/rectangles.html